OpenStudy (anonymous):

Help solve for "x" please: 2x - (5x - 4) = 4 (-x + 1)

OpenStudy (anonymous):

2x−(5x−4)=4(−x+1) 2x−5x+4=4(−x+1) −3x+4=4(−x+1) x=0

OpenStudy (anonymous):

Good luck

OpenStudy (anonymous):

I'm stuck with what to do at −3x+4=4(−x+1).

OpenStudy (anonymous):

−3x+4=4(−x+1) −3x+4=−4x+4 x=0

OpenStudy (anonymous):

Okay, thanks!